what is the ph of a 0.210 m solution of acetic acid? ka= 1.75x10 -5

1
Expert's answer
2022-03-29T13:53:06-0400

"pH = pKa + log(1)"

PH= 1.75× 10^-5 + log(0.210)

pH= 1.75+ 0.68= 2.43